a large group of friends decide to go to a football match. they fill 6 taxis and four go on foot. on the way back they fill four taxis and sixteen of them walk home.
make up an equation and solve it to find the number of people in each taxi

Answer by TimothyLamb(4379)

x = number of people per taxi
y = total size of group
---
6x + 4 = y
4x + 16 = y
---
put the system of linear equations into standard form
---
6x - y = -4
4x - y = -16
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of people per taxi = 6
y = total size of group = 40
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---